solvee the system by using the addition method
10x-5y=10 9x-6y=24 (__,__)

Answer by mananth(12270)

10x-5y=10
9x-6y=24
-6(10x-5y)+5(9x-6y)=-60+120
-60x+30y+45x-30y=+60
-15x=60
x=-4
Plug the value of x in 10x-5y=10
10*-4-5y=10
-5y=50
y=-10
Ananth